# VEAttack: Downstream-agnostic Vision Encoder Attack against Large Vision Language Models

### v1

@@ Abstract
Large Vision-Language Models (LVLMs) have demonstrated remarkable capabilities in multimodal understanding and generation, yet their vulnerability to adversarial attacks raises significant robustness concerns. While existing effective attacks always focus on task-specific white-box settings, these approaches are limited in the context of LVLMs, which are designed for diverse downstream tasks and require expensive full-model gradient computations. Motivated by the pivotal role and wide adoption of the vision encoder in LVLMs, we propose a simple yet effective Vision Encoder Attack (VEAttack), which targets the vision encoder of LVLMs only. Specifically, we propose to generate adversarial examples by minimizing the cosine similarity between the clean and perturbed visual features, without accessing the following large language models, task information, and labels. It significantly reduces the computational overhead while eliminating the task and label dependence of traditional white-box attacks in LVLMs. To make this simple attack effective, we propose to perturb images by optimizing image tokens instead of the classification token. We provide both empirical and theoretical evidence that VEAttack can easily generalize to various tasks. VEAttack has achieved a performance degradation of $94.5\%$ on image caption task and $75.7\%$ on visual question answering task. We also reveal some key observations to provide insights into LVLM attack/defense: 1) hidden layer variations of LLM, 2) token attention differential, 3) Möbius band in transfer attack, 4) low sensitivity to attack steps. The code is available at https://github.com/hfmei/VEAttack-LVLM
@@ 1 Introduction
Large Vision-Language Models (LVLMs) liu2023visual ; zhu2023minigpt ; Qwen-VL ; team2023gemini ; sun2021ernie have revolutionized multimodal AI by unifying pre-trained vision encoders radford2021learning ; li2022blip ; li2023blip with large language models (LLMs) touvron2023llama ; achiam2023gpt , enabling seamless integration of visual and textual understanding for tasks including visual question answering, image captioning, etc. However, their reliance on vision inputs inherits adversarial vulnerabilities well studied in computer vision madry2017towards ; szegedy2013intriguing , where imperceptible perturbations can mislead model predictions. As LVLMs increasingly deploy in real-world systems, their robustness against such attacks becomes critical, especially since adversarial perturbations to vision inputs can propagate through cross-modal alignment, causing catastrophic failures in downstream text generation schlarmann2023adversarial ; zhao2023evaluating ; bhagwatkar2024towards ; xu2024shadowcast ; yin2023vlattack .
Existing white-box attacks croce2020reliable ; schlarmann2023adversarial ; cui2024robustness ; bhagwatkar2024towards could effectively fool LVLMs by leveraging full-model gradients, such as APGD croce2020reliable , as shown in Figure 1. However, they face two key limitations: (1) Task and label specificity. LVLMs are inherently multi-task systems, but traditional white-box attacks optimize perturbations for a single task with specific labels (e.g., captioning), which makes the attacking cost proportional to the number of tasks. Although transferring generated adversarial examples to other tasks could be a naive solution to this issue, the nature of white-box attacks, i.e., task and label specificity, makes them ineffective when transferred to other tasks (e.g., VQA) according to our empirical evidence in Figure 2. (2) High computational costs. Even for single-task attacks, white-box methods incur significant computational overhead due to the growing data volumes and model parameters in LVLMs, which hinders their practicality for real-world applications yaras2024compressible ; wang2022efficientvlm ; liu2024minicache . A naive solution could be noise injection without access to the LVLM gradients, such as Gaussian noise. However, due to a lack of targeted optimization, such weak perturbations fail to mislead LVLM outputs effectively, as shown in Figure 1.
@@
Figure 1: The illustration of different attack paradigms where the white modules are accessible to the attacker, while the dark gray modules are inaccessible during the attack.
In this paper, we aim to explore an efficient form of white-box attack against LVLMs with strong generalization that attacks multiple tasks with a single perturbed image. We attribute this objective to the trade-off in the degree of reliance on task-specific gradients, where reducing such dependence seeks to improve generalization across multiple tasks and decrease computational costs, while maintaining effective adversarial perturbations. Drawing from traditional vision tasks, where a strong vision backbone enhances downstream task success lin2017feature ; liu2021swin ; tan2019efficientnet ; he2017mask ; goldblum2023battle , and recognizing the critical role of the vision encoder in LVLMs for overall performance shen2021much ; jain2024vcoder ; tong2024eyes ; yin2024survey , we target the vision encoder of LVLMs for attack. The Vision Encoder Attack (VEAttack) exploits the shared representations of LVLMs across tasks, enhancing generalization and improving computational efficiency. Specifically, we propose to minimize the cosine similarity between clean and perturbed visual features of the vision encoder. Our theoretical analysis reveals that such perturbations guarantee a lower-bound perturbation in the aligned features for LLMs, enabling effective adversarial attacks without task-specific dependencies. Although targeting the class token of the vision encoder could be a naive solution, our theoretical findings show that such perturbations have a reduced impact on aligned features compared to directly targeting image token features. Thus, we propose to optimize perturbations on image token features for stronger effects.
@@
Through focusing merely on the vision encoder, VEAttack reduces the parameter space for attack, achieving an 8-fold reduction in time costs compared to ensemble attack schlarmann2023adversarial . It also enables downstream-agnostic generalization across diverse tasks, achieving a performance degradation of 94.5% on the image caption task and 75.7% on the visual question answering (VQA) task. Furthermore, drawing upon extensive experiments and empirical analyses employing VEAttack, we reveal four key observations to provide insights into LVLM attack or defense: 1) hidden layer variations of LLM, 2) token attention differential, 3) Möbius band in transfer attack, 4) low sensitivity to attack steps. For example, VEAttack reveals strong cross-model transferability, where perturbations crafted for one vision encoder (e.g., CLIP radford2021learning ) compromise others (e.g., Qwen-VL Qwen-VL ), and uncovers a Möbius band phenomenon: enhancing vision encoder robustness bolsters LVLM robustness, yet paradoxically, adversarial samples targeting such robust encoders exhibit greater attack transferability across diverse LVLMs. The insights from VEAttack will inspire the research community to deepen exploration of LVLM vulnerabilities, fostering advancements in both attack and defense mechanisms to enhance the robustness and security of multimodal AI systems.
@@ 2 Related work
Adversarial robustness of LVLMs. Adversarial training (AT) rice2020overfitting ; zhang2019theoretically ; dong2023adversarial ; lin2024adversarial has proven effective in enhancing model robustness against adversarial attacks typically in image classification tasks. With the advent of vision-language models like CLIP radford2021learning , research has shifted toward ensuring robustness and generalization in these multimodal systems mao2022understanding ; wang2024pre ; yu2024text ; choi2024adversarial . FARE schlarmann2024robust improves the robustness of downstream LVLMs through an unsupervised adversarial fine-tuning scheme on the vision encoder. Recent studies hossain2024sim ; malik2025robust further advance LVLM robustness by optimizing adversarial fine-tuning strategies and scaling up the vision encoder wang2024revisiting . These works reveal the critical role of a robust vision encoder in LVLM security.
Adversarial attacks on LVLMs. Adversarial attacks on LVLMs have gained significant attention due to their vulnerability to imperceptible perturbations. Black-box attacks zhao2023evaluating ; zhang2024anyattack ; xu2024highly ; yin2023vlattack leverage self-supervised learning, diffusion models, or multimodal transferability to craft effective attacks. However, these methods require carefully designed transfer strategies tailored to LVLMs and computationally expensive gradient estimation. White-box attacks can adapt traditional attack methods madry2017towards ; croce2020reliable ; carlini2017towards to effectively target LVLMs, achieving high attack success rates schlarmann2023adversarial ; cui2024robustness ; bhagwatkar2024towards , for example, attack schlarmann2023adversarial adapts APGD croce2020reliable to targeted and ensemble attack. However, their task-specific strategies could lead to low generalization for multi-task system LVLMs. This work aims to address the trade-off between the reliance of white-box attacks on LVLMs and their attack effectiveness, proposing a general and efficient form of white-box attack.
@@ 3 Preliminary
An LVLM generally comprises a pre-trained vision encoder (e.g., CLIP radford2021learning ), a large language model (LLM) brown2020language ; touvron2023llama ; achiam2023gpt , and a cross-modal alignment mechanism to achieve joint understanding of images and text. As for the vision encoder CLIP $f_{CLIP}$, it typically generates two outputs, which are class token embedding $z_{cls}\in\mathbb{R}^{1\times{d_{cls}}}$ and image tokens $z_{v}\in\mathbb{R}^{n_{v}\times d_{v}}$. Formally, LVLMs process an image-text pair $((v,t)\in\mathcal{V}\times\mathcal{T})$ to generate a textual output conditioned on multimodal representations, where $\mathcal{V}$ represents the visual input space and $\mathcal{T}$ denotes the textual input space. In the process of prediction by LVLMs, we denote the vision encoder as $f_{V}$, which follows the relationship as $f_{CLIP}=\{z_{cls};f_{V}(v)\}$, $f_{V}(v)=z_{v}\in\mathbb{R}^{n_{v}\times d_{v}}$, where $n_{v}$ is the number of image tokens and $d_{v}$ is the dimension of each token. Then, the cross-modal alignment mechanism $f_{A}:\mathbb{R}^{n_{v}\times d_{v}}\to\mathbb{R}^{n_{v}\times d_{m}}$ aligns visual features $z_{v}$ with the input space of the language model, resulting in $z_{m}$. This feature will be combined with the textual tokens $z_{t}\in\mathbb{R}^{n_{t}\times d_{t}}$ from the tokenizer to form a unified input sequence. Finally, the LLM processes the input autoregressively to generate an output sequence, modeled as:
$p(y|z_{m};z_{t};\theta)=\prod_{i=1}^{N}p(y_{i}|y_{<i},z_{m};z_{t};\theta),$ (1)
where $i$ is the index of the current token position, $z_{m}=f_{A}(z_{v})=f_{A}(f_{V}(v))$, $z_{t}=tokenizer(t)$. In the traditional white-box adversarial attack croce2020reliable , the objective is to maximize the cross-entropy loss of the model on the adversarial input, which can be formalized as:
@@ 4 Methodology
In this section, we explore the vision encoder attack on LVLMs in a downstream-agnostic context. Specifically, we first analyze the challenges associated with traditional white-box attacks and ensure a lower-bound perturbation for aligned features in LVLMs. With a theoretical analysis of naive solutions within our framework, we redirect the attack target and achieve an effective VEAttack.
@@ 4.1 Redefine adversarial objective
To clarify our redefined adversarial objective, which shifts from full access attacks on LVLMs to targeting the vision encoder, we first establish definitions for different attacks. In traditional white-box attacks, attackers have full access to the LVLM’s architecture, including input image $v$, corresponding textual construction $t$, and model weights $\theta=\{\theta_{V},\theta_{A},\theta_{LLM}\}$, where the items in $\theta$ represent model weights for vision encoder $f_{CLIP}$, multimodal alignment mechanism $f_{A}$ and LLM $f_{LLM}$. In our VEAttack, we relieve the reliance of white-box setting, which transforms the full model attack into a vision encoder $f_{CLIP}$, where only model weights $\theta_{V}$ and input images $v$ are accessible. For an image $v$ with multi-task $T=\{T^{1},\cdots,T^{m}\}$, the corresponding constructions can be $t^{T}=\{t^{T^{1}},\cdots,t^{T^{m}}\}$. Following Eq. (2), the objective of the traditional white-box for multi-task can be formulated as:
@@
Figure 2: Comparison of transfer attack capability and time consumption between APGD and vision encoder attack (VEAttack).
Although traditional white-box attacks demonstrate potential adversarial capabilities schlarmann2023adversarial ; cui2024robustness ; bhagwatkar2024towards on a specific task $T^{i}$, it’s hard to maintain strong adversarial effectiveness when transferring attacks across different tasks. In Figure 2 (b), we conduct a transfer attack on two tasks $\{T^{1},T^{2}\}$, which are image captioning and visual question answering. For white-box APGD attack croce2020reliable , the performance degradation of adversarial sample $\tilde{v}^{T^{i}}$ on task $T^{j}$ is not obvious, where $i,j\in\{1,2\}$, $i\neq j$. Furthermore, it is also difficult for APGD to transfer to different downstream LLMs, like OpenFlamingo-9B to LLaVA-7B. Even in task-specific scenarios, white-box attacks typically entail higher computational costs. As shown in Figure 2 (a), the time cost of white-box attacks is about three times that of solely attacking the vision encoder on each specific task $T^{i}$. To address the above challenges, we propose targeting the vision encoder to craft adversarial perturbations $\tilde{v}$ that exploit its shared representations across tasks $T^{i}$, leveraging its critical role in LVLMs shen2021much ; jain2024vcoder ; tong2024eyes ; yin2024survey and insights from strong vision backbones in traditional tasks lin2017feature ; liu2021swin ; tan2019efficientnet .
@@
For LLaVa liu2023visual with a linear alignment layer, let $\Delta z_{v}=\tilde{z}_{v}-z_{v}$ denote the difference between the image tokens output by the vision encoder CLIP before and after the perturbation, $\|\Delta z_{v}\|_{F}\geq\Delta$, $W_{a}$ is the weight of projection layer, $\sigma_{min}$ denotes the minimum singular value of $W_{a}$. Assume $\sigma_{min}(W_{a})>0$, then the difference between aligned features $z_{m}=f_{A}(z_{v})$ and $\tilde{z}_{m}=f_{A}(\tilde{z}_{v})$ for downstream LLMs will satisfy $\|\Delta z_{m}\|_{F}\geq\sigma_{min}(W_{a})\Delta>0$.
Proof. See Appendix A.
Figure 3: The feature difference before and after VEAttack with different perturbation budgets.
The Proposition 1 gives a lower bound of perturbations on the aligned feature $z_{m}$ when attacking on $z_{v}$ of the vision encoder, so the perturbations from the VEAttack on $z_{v}$ can reliably influence downstream $f_{LLM}$, ensuring the feasibility of VEAttack. When we attack the vision encoder of LVLMs, the perturbed feature $\tilde{z}_{v}$ will be represented as a fundamental feature to affect any downstream tasks and LLMs. In Figure 3, we empirically show that perturbations in the vision encoder features $\tilde{z}_{v}$ propagate to the aligned features $\tilde{z}_{m}$, with increasing perturbation amplitude amplifying the feature differences in both $\Delta z_{v}$ and $\Delta z_{m}$. Following the important role of the vision encoder and theoretical analysis of the propagability of perturbations, we redefine the adversarial objective to the vision encoder and reformulate the attack objective in Eq. (3) as follows:

### v2

@@ Abstract
Large Vision-Language Models (LVLMs) have demonstrated capabilities in multimodal understanding, yet their vulnerability to adversarial attacks raises significant concerns. To achieve practical attacking, this paper aims at efficient and transferable untargeted attacks under limited perturbation sizes. Considering this objective, white‑box attacks require full‑model gradients and task‑specific labels, making costs scale with tasks, while black‑box attacks rely on proxy models, typically requiring large perturbation sizes and elaborate transfer strategies. Given the centrality and widespread reuse of the vision encoder in LVLMs, we adopt a gray‑box setting that targets the vision encoder alone for efficient but effective attacking. We theoretically establish the feasibility of vision‑encoder‑only attacks, laying the foundation for our gray‑box setting. Based on this analysis, we propose perturbing patch tokens rather than the class token, informed by both theoretical and empirical insights. We generate adversarial examples by minimizing the cosine similarity between clean and perturbed visual features, without accessing the subsequent models, tasks, or labels. This significantly reduces computational overhead while eliminating the task and label dependence. VEAttack has achieved a performance degradation of $94.5\%$ on image caption task and $75.7\%$ on visual question answering task. We also reveal some key observations to provide insights into LVLM attack/defense: 1) hidden layer variations of LLM, 2) token attention differential, 3) Möbius band in transfer attack, 4) low sensitivity to attack steps. The code is available at https://github.com/hefeimei06/VEAttack-LVLM.
@@ 1 Introduction
Large Vision-Language Models (LVLMs) (Liu et al., 2023; Zhu et al., 2023; Bai et al., 2023; Team et al., 2023) have revolutionized multimodal AI by unifying pre-trained vision encoders (Radford et al., 2021; Li et al., 2022) with large language models (LLMs) (Touvron et al., 2023; Achiam et al., 2023), enabling seamless integration of visual and textual understanding for tasks including visual question answering, image captioning, etc. However, their reliance on vision inputs inherits adversarial vulnerabilities well studied in computer vision (Madry et al., 2018; Szegedy et al., 2014), where imperceptible perturbations can mislead model predictions. As LVLMs increasingly deploy in real-world systems, their robustness against such attacks becomes critical, especially since adversarial perturbations to vision inputs can propagate through cross-modal alignment, causing catastrophic failures in downstream text generation (Zhao et al., 2023; Cui et al., 2024; Li et al., 2025).
We explore the vulnerability of LVLMs under untargeted attacks with small perturbation sizes, which follows the standard attack setting (Madry et al., 2018). However, since LVLM targets different tasks with high computational cost, we aim at efficient attacks that can transfer across models and tasks to ensure practical attacks, which existing attacks can hardly achieve. Specifically, white-box attacks (Schlarmann and Hein, 2023; Bhagwatkar et al., 2024) normally optimize perturbations for a single task with specific labels (e.g., captioning), making the cost proportional to the number of tasks. Although transferring adversarial examples to other tasks could be a naive solution, the nature of white-box attacks, i.e., task and label specificity, makes them ineffective when transferred to other tasks (e.g., VQA) according to our empirical evidence in Fig. 2. Black-box attacks (Zhang et al., 2024; Jia et al., 2025) typically manipulate LVLMs to output specified targets with proxy known vision encoders (e.g., CLIP-ViT-B/16) and substantial perturbation budgets. As shown in Fig. 10 and Fig.2 (b), when confronted with minimal perturbations, these methods often struggle to achieve effective attacks. One potential solution is the gray-box setting in Fig. 1, where partial model parameters are accessible, which reduces the attacking cost compared to white-box while enhancing attacking capability compared to black-box. However, it is difficult for existing gray-box attacks (Wang et al., 2024c) to achieve better satisfactory performance, even with text modality information and additional text encoders. Drawing from traditional vision tasks, where strong vision backbones boost downstream performance (Lin et al., 2017; Liu et al., 2021), and the pivotal role of vision encoders in LVLMs (Jain et al., 2024; Tong et al., 2024), we argue that the attacking capability of the vision encoder has not been fully explored and propose to investigate the feasibility of leveraging the vision encoder solely to achieve efficient and transferable attacks on LVLMs.
In this paper, we attribute this objective to explore the attack potential of the vision encoder and ensure the effectiveness of perturbation propagation to downstream LLMs, so that vision-encoder-only attack could improve generalization across multiple tasks and decrease computational costs, while maintaining effective adversarial perturbations. Specifically, we theoretically analyze the lower bound of perturbation in the multimodal aligned features for LLMs when perturbing solely on vision encoders, which provides a feasibility foundation for our attack context and gray-box attacks. Through proposing naive solutions to our question drawn from black-box attack (Zhao et al., 2023) and adversarial finetuning work (Schlarmann et al., 2024), our theoretical findings show that such perturbations have a reduced impact on class token features compared to directly targeting patch token features. Thus, we optimize perturbations of vision encoders by minimizing the cosine similarity between clean and perturbed patch token features for stronger attack effects. Through focusing merely on the vision encoder, VEAttack reduces the parameter space for attack, achieving an 8-fold reduction in time costs compared to ensemble attack (Schlarmann and Hein, 2023). It also enables downstream-agnostic generalization across diverse tasks, achieving a performance degradation of 94.5% on the image caption task and 75.7% on the visual question answering (VQA) task.
@@
Figure 1: The illustration of different attack paradigms where the white modules are accessible to the attacker, while dark modules are inaccessible during the attack with $2/255$ as perturbation budgets.
@@
Furthermore, drawing upon extensive experiments and empirical analyses employing VEAttack, we reveal four observations to provide insights into LVLM attack or defense: 1) hidden layer variations of LLM, 2) token attention differential, 3) Möbius band in transfer attack, 4) low sensitivity to attack steps. For example, VEAttack reveals strong cross-model transferability, where perturbations crafted for one vision encoder (e.g., CLIP (Radford et al., 2021)) compromise others (e.g., Qwen-VL (Bai et al., 2023)) with a Möbius band phenomenon: enhancing vision encoder robustness bolsters LVLM robustness, yet paradoxically, adversarial samples targeting such robust encoders exhibit greater attack transferability across diverse LVLMs. The insights from VEAttack will inspire the research community to deepen exploration of LVLM vulnerabilities, fostering advancements in both attack and defense mechanisms to enhance the robustness and security of multimodal AI systems.
@@ 2 Related work
Adversarial robustness of LVLMs. Adversarial training (AT) (Rice et al., 2020; Dong and Xu, 2023; Lin et al., 2024) has proven effective in enhancing model robustness against adversarial attacks, typically in image classification tasks. With the advent of vision-language models like CLIP (Radford et al., 2021), research has shifted toward ensuring robustness and generalization in these multimodal systems (Mao et al., 2023; Wang et al., 2024b). FARE (Schlarmann et al., 2024) improves the robustness of downstream LVLMs through an unsupervised adversarial fine-tuning scheme on the vision encoder. Recent studies (Hossain and Imteaj, 2024; Malik et al., 2025) further advance LVLM robustness by optimizing adversarial fine-tuning strategies and scaling up the vision encoder. These works reveal the critical role of a robust vision encoder in LVLM security.
Adversarial attacks on LVLMs. Adversarial attacks on LVLMs are normally divided into white-box, black-box, and gray-box attacks. White-box attacks (Schlarmann and Hein, 2023; Cui et al., 2024) can adapt traditional attack methods (Madry et al., 2018; Croce and Hein, 2020) to effectively target LVLMs, achieving high attack success rates. However, their task-specific strategies could lead to low generalization for multi-task system LVLMs. CroPA (Luo et al., 2024) proposes learnable prompts to effectively mislead VLMs across diverse textual inputs, but requires calculating gradients through the massive LLM. Black-box attacks (Zhao et al., 2023; Zhang et al., 2024; Li et al., 2025; Jia et al., 2025) leverage self-supervised learning, diffusion models, or other transfer strategies to craft effective attacks with relatively larger perturbations. Doubly-UAP (Kim et al., 2024) introduced universal perturbations trained on large-scale datasets to achieve cross-image transferability with significant pre-training costs. Gray-box attacks (Wang et al., 2024c; a; Zhang et al., 2025; Wang et al., 2023) relieve the attack objective to the visual features in LVLMs, but still introduce text information and the additional text encoder for better performance. Different from the gray-box context above, our work aims to explore using only the vision encoder and image modality to achieve a more effective and efficient attack.
@@ 3 Preliminary
An LVLM generally comprises a pre-trained vision encoder (e.g., CLIP (Radford et al., 2021)), a large language model (LLM) (Brown et al., 2020; Touvron et al., 2023; Achiam et al., 2023), and a cross-modal alignment mechanism to achieve joint understanding of images and text. As for the vision encoder CLIP $f_{CLIP}$, it typically generates two outputs, which are class token embedding $z_{cls}\in\mathbb{R}^{1\times{d_{cls}}}$ and patch tokens $z_{v}\in\mathbb{R}^{n_{v}\times d_{v}}$. Formally, LVLMs process an image-text pair $((v,t)\in\mathcal{V}\times\mathcal{T})$ to generate outputs conditioned on multimodal representations, where $\mathcal{V}$ represents the visual input space and $\mathcal{T}$ denotes the textual input space. In the process of prediction by LVLMs, we denote the vision encoder as $f_{V}$, which follows the relationship as $f_{CLIP}=\{z_{cls};f_{V}(v)\}$, $f_{V}(v)=z_{v}\in\mathbb{R}^{n_{v}\times d_{v}}$, where $n_{v}$ is the number of patch tokens and $d_{v}$ is the dimension of each token. Then, the cross-modal alignment mechanism $f_{A}:\mathbb{R}^{n_{v}\times d_{v}}\to\mathbb{R}^{n_{v}\times d_{m}}$ aligns visual features $z_{v}$ with the input space of the language model, resulting in $z_{m}$. This feature will be combined with the textual tokens $z_{t}\in\mathbb{R}^{n_{t}\times d_{t}}$ from the tokenizer to form a unified input sequence. Finally, the LLM processes the input autoregressively to generate an output sequence, modeled as:
$\vskip-5.0ptp(y|z_{m};z_{t};\theta)=\prod_{i=1}^{N}p(y_{i}|y_{<i},z_{m};z_{t};\theta),$ (1)
where $i$ is the index of the current token position, $z_{m}=f_{A}(z_{v})=f_{A}(f_{V}(v))$, $z_{t}=tokenizer(t)$. In the traditional white-box adversarial attack (Croce and Hein, 2020), the objective is to maximize the cross-entropy loss of the model on the adversarial input, which can be formalized as:
@@ 4 Methodology
In this section, we explore the gray-box attack on LVLMs in a vision-encoder-only context. Specifically, we first analyze the challenges associated with existing white-box and black-box attacks under our gray-box settings, while ensuring a lower-bound perturbation for aligned features in LVLMs as a feasibility basis. With a theoretical analysis of naive solutions within our framework, we redirect the attack target and achieve an effective VEAttack.
@@ 4.1 Redefine adversarial objective
To clarify our redefined adversarial objective, which shifts from full access attacks on LVLMs to targeting the vision encoder, we first establish definitions for different attacks. In traditional white-box attacks, attackers have full access to the LVLM’s architecture, including input image $v$, corresponding textual construction $t$, and model weights $\theta=\{\theta_{V},\theta_{A},\theta_{LLM}\}$, where the items in $\theta$ represent model weights for vision encoder $f_{CLIP}$, multimodal alignment mechanism $f_{A}$ and LLM $f_{LLM}$. In our VEAttack, we relieve the reliance on the white-box setting, which transforms the full model attack into a vision encoder $f_{CLIP}$, where only model weights $\theta_{V}$ and input images $v$ are accessible. For an image $v$ with multi-task $T=\{T^{1},\cdots,T^{m}\}$, the corresponding constructions can be $t^{T}=\{t^{T^{1}},\cdots,t^{T^{m}}\}$. Following Eq. (2), the objective of the traditional white-box for multi-task can be formulated as:
@@
Figure 2: Transfer capability and time consumption of white-box APGD, black-box FOA-Attack, and our gray-box VEAttack.
Although white-box attacks demonstrate potential adversarial capabilities (Schlarmann and Hein, 2023; Cui et al., 2024) on a specific task $T^{i}$, it’s hard to maintain strong adversarial effectiveness when transferring attacks across different tasks. Black-box attacks (Li et al., 2025; Jia et al., 2025) could effectively enhance the generalization but need time-consuming transfer strategies and larger perturbations as shown in Fig. 10. In Fig. 2 (b), we conduct a transfer attack on two tasks $\{T^{1},T^{2}\}$, which are image captioning and visual question answering with perturbation budget $\epsilon=4/255$. For white-box APGD (Croce and Hein, 2020) and black-box FOA-Attack (Jia et al., 2025), the performance degradation of adversarial sample $\tilde{v}^{T^{i}}$ on task $T^{j}$ is not obvious, where $\in\{1,2\}, iji\neq j$. Even in task-specific scenarios, white-box and black-box attacks typically entail higher computational costs as shown in Fig. 2 (a). Facing the above challenges, we propose targeting the vision encoder to craft adversarial perturbations $\tilde{v}$ that exploit its shared representations across tasks $T^{i}$, leveraging its critical role in LVLMs (Shen et al., 2021; Tong et al., 2024) and insights from vision backbones in traditional tasks (Lin et al., 2017; Liu et al., 2021).
@@
For LLaVa (Liu et al., 2023) with a linear alignment layer, let $\Delta z_{v}=\tilde{z}_{v}-z_{v}$ denote the difference between the patch tokens output by the vision encoder CLIP before and after the perturbation, $\|\Delta z_{v}\|_{F}\geq\Delta$, $W_{a}$ is the weight of projection layer, $\sigma_{min}$ denotes the minimum singular value of $W_{a}$. Assume $\sigma_{min}(W_{a})>0$, then the difference between aligned features $z_{m}=f_{A}(z_{v})$ and $\tilde{z}_{m}=f_{A}(\tilde{z}_{v})$ for downstream LLMs will satisfy $\|\Delta z_{m}\|_{F}\geq\sigma_{min}(W_{a})\Delta>0$.
Proof. See Appendix B.
Figure 3: The feature difference before and after VEAttack with different budgets.
The Proposition 1 gives a lower bound of perturbations on the aligned feature $z_{m}$ when attacking on $z_{v}$ of the vision encoder, so the perturbations from the VEAttack on $z_{v}$ can reliably influence downstream $f_{LLM}$, ensuring the feasibility of VEAttack. When attacking the vision encoder of LVLMs, the perturbed feature $\tilde{z}_{v}$ will be represented as a fundamental feature to affect any downstream tasks and LLMs. In Fig. 3, we empirically show that perturbations in the features $\tilde{z}_{v}$ propagate to the aligned features $\tilde{z}_{m}$, with increasing perturbation amplitude amplifying the feature differences in both $\Delta z_{v}$ and $\Delta z_{m}$. Following the importance of the vision encoder and theoretical analysis of the propagability of perturbations, we redefine the adversarial objective to the vision encoder and reformulate the attack objective in Eq. (3) as follows: